OpenStudy (anonymous):

Find the inverse of the function. Enter NO INVERSE if no inverse exists. f (x) = 8x2 + 9

OpenStudy (anonymous):

x=8y^2+9 x-9=8y^2 (x-9)/8=y^2 square root ((x-9)/8)=y

OpenStudy (anonymous):

the answer is NO INVERSE

OpenStudy (anonymous):

Oh.. then I don't know sorry :(

OpenStudy (anonymous):

Well you have to specify the domain!

OpenStudy (anonymous):

Below is quotation from Purple Math site http://www.purplemath.com/modules/invrsfcn2.htm: "Since the original function had two points that shared the same Y-VALUE, then the inverse of the original function will not be a function. This means, for instance, that no parabola (quadratic function) will have an inverse that is also a function."